A construction company builds a rectangular fenced in area with a perimeter of 46 meters. The length is 5 meters more than the width. what are the dimensions of the fence?
Answer by TimothyLamb(4379)

---
2x + 2y = 46
x = y + 5
---
put the system into standard form:
---
2x + 2y = 46
x - y = 5
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
answer:
L= 14
w= 9
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php